Calculate voltage and current across RL using mesh or nodal analysis. Verify using Thevenin’s and Norton’s Theorems. (RL = 1.5k)3. 13 3.3ke3 Prelab 5 Task li Calculate voltage across R and current using mesh or nodal analysis. Verify using Thevenins & N
Answer #1

In Mesh - . KULegn 5 + 2.78 103 (I, Ig) +3.3 x 103 (I – I3)=0 6 I, - d. 7 Ig -3.3 Iz = 15* 4.5V R 1 Re 4 -13)=0 ISUT O H 2 -3Trika $ 3.3KA using The venins Theoremy for equivalent Resistance Rth (Thevenins Resistance) voltage source & shortckt - 66From EQ n f . 61 – 6 Ig = 15x10 3 -644 +12 19 = 0 I GI2 = 15x103 Is = d.5mA Ij = 5 MA Now for Uth in mesh 6 → 8.3 x 103 I – 3